OpenStudy (anonymous):

A number added to 8 times that numbers recipricol is 6, find the number a 3 b 2 c 7 d 14

OpenStudy (anonymous):

Call your unknown number x. The reciprocal would be \[\frac{1}{x}\]You are supposed to add your unknown number to 8 times its reciprocal, or \[x+8(\frac{1}{x})\]This is all supposed to equal 6, so you have \[x+\frac{8}{x}=6\] Solve for x to get your number.

OpenStudy (anonymous):

n + 8(1/n) = 6 n^2 + 8 = 6n n^2 - 6n + 8 = 0 (n - 4)(n - 2) = 0 n = 2, 4

OpenStudy (anonymous):

is 2 its answer ??????

OpenStudy (anonymous):

Out of the choices given, yes. However, 4 is also a solution.

OpenStudy (anonymous):

2 and 4 are answers thank you

OpenStudy (anonymous):

L^2+8=6L L^2-6L+8=0 L=2,4